OpenStudy (anonymous):

How do I write this in interval notation: {x element R : x>=14}

OpenStudy (anonymous):

\([14,\infty)\)

OpenStudy (anonymous):

\[ \geq 14\to [4, \]

OpenStudy (anonymous):

No bounds on how big so:\[ \to \infty) \]

OpenStudy (anonymous):

Thank you both very much.